# Concerns and Values in Human-Robot Interactions: A Focus on Social Robotics

### v1

@@ 3.2.1 Application topics
The paper’s application topics were extracted with the thematic analysis approach [13] as in previous works [91, 21], resulting in 10 themes areas emerged from the survey papers, reported in Figure 3. 24 papers reported works on healthcare (e.g., elderly healthcare [5]), 6 on ethics (e.g., [19]), 5 on education (e.g., [97]), 2 on trust (e.g., [75]), 1 on feminism, 1 on manipulation, 1 on attachment, 1 on home, 1 on mental health, and 1 on human-robot relationships.
@@ 3.2.2 Cultural background
The country of affiliation of the authors of the publications are from North America and Western Europe. While the country of affiliation of the authors is not necessarily indicative of their cultural background, this information provides an idea of the point of view of the contributions included in the review, which is Western-centric. Of the 43 publications considered, 25 featured an empirical study. In these cases, knowing the population’s country of origin, or the place where the study was executed, is relevant. This geographical distribution mirrors the authors’ case, although it includes data from Brazil and China. Regarding the population’s age, the studies encompass young, adult and elderly people interacting with social robots, and in three cases they include other stakeholders, e.g., parents, teachers, and caretakers.
@@ 3.2.3 Transcripts
We report the initial topics and values in Appendix D while Tables 2 and 3 collect the papers included in the scoping review and the corresponding values and topic defined at the end of the design process are presented in detail in Section 4.3.
@@ 3.3 Preliminary discussion and limitations
Results from Phase 1 contribute to the field of HRI by offering an overview of the current ethical discussions around values in the context of healthcare, education, and domestic HRI. Within application contexts, healthcare emerged as the most represented domain, suggesting that ethical considerations are especially relevant in sensitive contexts where robots directly impact human health and wellbeing. Our results show that other topics like feminism and manipulation were largely underexplored. However, this could have been affected by the choice of the search query that only included domains of healthcare, education, and domestic. Future research could investigate these areas to provide a more holistic view of ethical considerations in HRI. The Western-centric perspective of the surveyed publications points to a significant gap in the cultural representation of values in HRI. Future work should explore more cultural diversity in different socio-cultural contexts.
@@ 4 Phase 2: Focus Groups
Building on the list of values that emerged from the literature review, Phase 2 aims to define a comprehensive list of topics and values specific to human-robot interaction (RQ2). To achieve this, we conducted two focus groups with six experts in technology ethics and philosophy. These experts participated in the focus groups consisting of three activities to develop a list of values in the context of HRI.
@@ 4.2 Focus group activities
We conducted the focus groups, which lasted about two hours, using an online video conference tool. To facilitate the activities, we used the online tool Miro [^3] https://miro.com/, which allows users to collaborate online using interactive boards that can be prepared in advance.
These focus groups consisted of three activities, namely (1) HRI scenarios analyses (individual task), (2) values review (group task), and (3) contextual information and relationship with values (group task). These activities are inspired by previous research in related fields. For example, Spitale et al. [89] utilised a scenario-based evaluation to assess whether the language generated by large language models was appropriate for a robotic coach promoting mental wellbeing. Similarly, activities involving reviewing existing findings, such as literature reviews, have been commonly employed in prior studies (e.g., [16]. Additionally, linking contextual information with the task at hand is a well-established practice in Interaction Design studies [40]. Two researchers guided and presented the activities.
@@ 4.2.1 Activity 1: scenarios analysis
The first activity involved analysing four distinct HRI fictional value scenarios provided to the participants. A value scenario is an extension of scenario-based design that helps envision the systemic effects of new technologies [64]. We developed our HRI value scenarios through a structured process. First, we identified the main topics for our value scenarios based on the results of the previous phase and previous research [53]. Specifically, we considered healthcare (focused both on elderly care [14] and autism therapy [28]), education (focused on children’s learning [63]), and home assistance [94] as our application contexts (matching the keywords of the search query in Phase 1). Second, we created four value scenarios, one for each identified HRI domain, using ChatGPT [^4] https://chatgpt.com/, as demonstrated by Jung et al. [50]. We followed the same approach to draft prompts for generating these scenarios. The specific prompts used to generate the scenarios are provided in Appendix A. Third, we refined the generated value scenarios by removing any references to potential moral values to avoid biasing the scenarios. The scenarios generated are reported in Appendix B.
We provided participants with these value scenarios and asked them to individually identify the moral values present in each scenario. This activity lasted approximately 40 minutes, 10 minutes for each scenario.
@@ 4.2.2 Activity 2: values review
The second activity involved presenting participants with the results from Phase 1, which included a list of topics and values identified and extracted from the literature. To make these HRI moral values easier to visualise, we organised them in both a table (see Table 6) and a wheel (as shown in Figure 5). We first asked participants to take their time reviewing the values, then encouraged them to collaboratively discuss and reflect on them. This activity lasted approximately 20 minutes.
@@ 4.2.3 Activity 3: contextual information and relationship with values
The final activity involved discussing the relationship between the identified values and the contextual information. Our literature review revealed that most values are highly context-dependent in line with previous literature [27]. Therefore, we decided to prioritise context as the main lens for framing these values more effectively in any HRI scenario. To this end, we have designed an activity where we asked participants to discuss the relationship between the values they have reviewed and the contextual information of HRI scenarios taking inspiration from theories in Contextual Design [45]. We specifically prompted them using the following contextual questions (as depicted in Figure 4): (1) Who are the users and robots? What are their demographics (age, gender, education, etc.)? What are their roles and responsibilities?; (2) What are the users’ goals? What are the primary tasks they need to accomplish? What are their motivations for using the robot?; (3) Where is the robot being used? Is it being used in a specific physical location (e.g., home, office, outdoors)? What environmental factors (lighting, noise, space) influence its use?; (4) When (how long) is the robot being used? Is there a specific time of day, week, or year when the robot is used most? How does the timing affect the user’s interaction with the robot?; (5) Why is the robot being used? What problems or needs does the robot address? What are the user’s expectations from the robot?; (6) How is the robot being used? What steps do users take to accomplish their tasks? What tools or devices do they use in conjunction with the robot?; (7) What are the social and cultural contexts? Are there social norms or cultural factors that influence how the robot is used? How do interactions with other people (colleagues, family, friends) affect usage?; (8) What are the constraints and limitations? Are there any physical, cognitive, or technological limitations users face? What budgetary, time, or resource constraints exist? This activity lasted around one hour.
@@ 4.2.4 Data analysis
We first automatically transcribed the two focus group recordings and then manually reviewed them for accuracy. We used the thematic analysis method [13], with the values identified during the scoping review serving as main themes and the topics as codes.
@@ 4.3 Results: topics and values in HRI
In Activity 1, the participants were asked to identify the values in the fictional HRI scenarios presented to them, without any influence from our preliminary scoping review findings. Activity 2 revealed the gaps in the values identified by the scoping review, while Activity 3 highlighted the link between contextual information and moral values. The participants noted that each piece of contextual information could be associated with multiple moral values, though they focused on the most relevant connections during discussions. At the end of the workshop, we revised the initial value definitions from the scoping review, and based on the outputs of the focus groups, we defined 26 topics grouped under 8 values, as shown in Table 4. Each topic has a positive or negative valence and is visualised in Figure 5. For clarity, we combined the results from all activities and we report the final list in the following sections. Appendix E provides a detailed overview of the thematic analysis, including citations, quotes, codes, and themes from participants organised by activity. Our analysis reached data saturation [44], allowing us to confidently identify and understand the final list of values, which are described in the following paragraphs. Note that this is a simplified and reductionist framework designed to assist researchers in identifying and addressing values within the HRI scenarios they are developing and not a comprehensive tool that captures the full complexity of ethical concerns and values, cultural contexts, and stakeholders’ perspectives in the HRI realm.
@@ 4.3.1 Value: Agency
This value encompasses three topics – independence, dependence, and control – all related to the user’s physical freedom. This term, prior to the focus groups, referred to one of the topics, but it was promoted to a value following the suggestions from the focus groups participants, uniforming our definition of agency with Prunkl [72]: “A person is able to act on the beliefs and values they hold. This implies that they have meaningful options available to them, allowing them to make choices that are of practical import to their life”. As it emerged during the focus groups, agency differs from autonomy as the latter is used in this work to encompass topics which do not possess physical connotations. The related topics are defined as follows.
@@
The use of robots in daily life can significantly enhance a user’s independence, by reducing their physical dependence on other people, contributing to a sense of personal agency. This is especially evident in the case of elderly individuals or people with physical disabilities who require constant assistance. We included in this topic the topics about ease of use of the robot appliance, which should not require the user to rely on someone else to use the robot, as emerged during the scoping review [93, 99].
@@
The previous topic can also be viewed with negative connotations [19, 17]. Indeed, when the subject’s independence relies on robots instead of other humans, the subject becomes dependent on the robots themselves. For example, the focus groups’ participants mentioned that this could lead to a deterioration of various skills, such as cooking or reasoning. One of the participants particularly questioned if deskilling might be offset by the development of new, possibly superior, skills gained through interaction with the robot.
@@
To ensure the user’s agency, robots should be designed to never physically constrain or force the user into actions they do not wish to take. During the scoping review, this topic was initially categorised under the value empowerment (e.g., [99]). However, experts suggested removing empowerment as a separate value and instead placing control under the value agency. The required level of control can vary greatly depending on the context in which a robot is used, as noted in the focus groups discussions. For example, a robot assisting in physical rehabilitation should follow the user’s pace and respond to their commands rather than pushing them too hard or acting autonomously in a way that feels coercive.
@@ 4.3.2 Value: Connectedness
This value covers the social aspects involved in human-robot interactions, especially in older adults [73]. The focus groups participants suggested that this moral principle is linked to whether the robot is replacing a human role, enhancing connections, or enabling new activities. They emphasised that the robot should not replace humans, but support them.
@@
Robots have the potential to provide valuable companionship for individuals who feel isolated or lonely, e.g., [51, 46]. A companion robot can engage in conversations, respond to emotional cues, and even provide comfort just through its presence. For example, elderly individuals who live alone might find a robotic companion to be a source of emotional support, helping to alleviate feelings of loneliness.
@@
Beyond providing direct companionship, robots can also facilitate human connections, as highlighted by Amirabdollahian et al. [6]. For example, robots can help the elderly stay in touch with their families. This can happen also when they serve as conversation starters or mediators in social settings during group events. In other instances, they achieve the same objective by connecting the user with friends and family through video calls.
@@
On the negative side, while robots can enhance social connectedness, there is also a risk that their use might become a substitute for human interaction as emerged during the scoping review (e.g., [109, 93]). This could happen when a person begins to prefer the company of their robot over that of other people. The focus groups participants provided several examples of this. For instance, family members might feel less compelled to check on their elderly relatives, knowing that a robot is providing companionship. Similarly, teachers may not engage effectively with students or provide meaningful feedback if a robot takes over all teaching activities.
@@ 4.3.3 Value: Privacy
The value of privacy [32] is the one encompassing the largest amount of topics as highlighted in both the scoping review (e.g., in [87]) and the focus groups. It includes – among others – the right to be informed, to access and share the data collected, and the issues related to being continuously under the watchful eye of a robot. The focus groups participants have also highlighted that this value is strongly related to where, how and in which context the robot is used. Using a robot at home is completely different from using it in a public space. The experts emphasised that privacy should be maintained even when users do not want others to know they are using the robot.
@@
From the scoping review emerged the importance of ensuring that the user has access to the data collected by the robot [30]. Additionally, it was noted that sharing recorded information with selected individuals can be beneficial, for instance, sharing biometric data with a chosen physician. The focus groups discussions emphasised that data protection is particularly crucial when dealing with vulnerable populations, such as children and the elderly. It is also vital to clearly disclose upfront who will have access to the collected data.
@@
One of the potential downsides of robots is the risk of invading the privacy of bystanders who have not consented to being monitored, e.g., [47]. This is a concern especially in hospitals, as the robot is surrounded by vulnerable subjects in delicate settings. For example, the focus groups participants questioned whether patients in nursing homes can truly give informed consent to data collection and use, given their potentially limited understanding.
@@
The constant presence of a robot in a user’s personal space can lead to feelings of discomfort and a sense of being constantly observed that can violate the value of intimacy, e.g., [4]. Furthermore, this can lead to an unconscious change in the users’ actions, behaving in ways they think will be acceptable to potential observers, rather than acting naturally. This resembles the effects of the panopticon [35], as noted by one of the focus groups participants. Even when the robot is not actively monitoring the environment, its presence alone is enough to influence the user’s behaviour.
@@
This topic covers all those concerns related to the excessive quantities of data collected by robotic devices (e.g., in [66, 46]). The continuous and aggressive collection of personal data for profiling risks misuse by both collecting companies and malicious actors. The focus groups participants expressed concerns that data from elderly care centres could be shared with insurance companies or sold to third parties without benefitting the user. These concerns are heightened for vulnerable groups like children, where parents or guardians should manage data consent. Additionally, the context of the application, such as in-home settings with less stringent regulations, can increase potential risks and harms, such as misuse by domestic abusers [107]. Other concerns, emerged during the scoping review, touch on the safe transmission and storage of this data.
@@
Transparency about the robot’s capabilities, especially regarding data collection and monitoring, emerges with positive connotations in the scoping review. Users should be fully informed about what the robot is capable of, and what data is being shared, as highlighted by focus groups participants. This can be tricky to implement when transparency would reduce the effectiveness of the robot itself, in the case of personal disclosure for instance.

### v2

@@ 3.2.1 Application topics
The paper’s application topics were extracted with the thematic analysis approach [braun2012thematic] as in previous works [spitale2023systematic, catania2023conversational], resulting in 10 areas, reported in Figure 3. 24 papers reported works on healthcare (e.g., elderly healthcare [amirabdollahian2013assistive]), 6 on ethics (e.g., [cappuccio2021can]), 5 on education (e.g., [van2020teachers]), 2 on trust (e.g., [rosero2023using]), 1 on feminism, 1 on manipulation, 1 on attachment, 1 on home, 1 on mental health, and 1 on human-robot relationships.
@@ 3.2.2 Cultural background
The country of affiliation of the authors of the publications is from North America and Western Europe. While the country of affiliation of the authors is not necessarily indicative of their cultural background, this information provides an idea of the point of view of the contributions included in the review, which is Western-centric. Of the 43 publications considered, 25 featured an empirical study. In these cases, knowing the population’s country of origin, or the place where the study was executed, is relevant. This geographical distribution mirrors the authors’ case, although it includes data from Brazil and China. Regarding the population’s age, the studies encompass young, adult and elderly people interacting with social robots, and in three cases they include other stakeholders, e.g., parents, teachers, and caretakers.
@@ 3.2.3 Transcripts
We report the initial topics and values in Table 2 while Tables 3 and 4 collect the papers included in the scoping review and the corresponding values and topic defined at the end of the design process are presented in detail in Section 4.3.
@@ 3.3 Preliminary discussion and limitations
Results from Phase 1 contribute to the field of HRI by offering an overview of the current ethical discussions around values in the context of healthcare, education, and domestic HRI. Within application contexts, healthcare emerged as the most represented domain, suggesting that ethical considerations are especially relevant in sensitive contexts where robots directly impact human health and wellbeing. Our results show that other topics like feminism and manipulation were largely underexplored. However, this could have been affected by the choice of the search query that only included domains of healthcare, education, and domestic. Future research could investigate these areas to provide a more holistic view of ethical considerations in HRI. The Western-centric perspective of the surveyed publications points to a significant gap in the cultural representation of values in HRI. Future work should explore more cultural diversity in different socio-cultural contexts. While we chose the interaction as a starting point for the value identification, the results are influenced by other elements, such as the stakeholders’ values and socio-cultural aspects, with effects on the robot’s role beyond the interaction itself.
@@ 4 Phase 2: Focus Groups
Building on the list of values that emerged from the literature review, Phase 2 aims to define a comprehensive list of topics and values specific to human-robot interaction (RQ2). To achieve this, we conducted two focus groups with six experts in technology ethics and philosophy, consisting of a discussion driven by three activities, to develop a list of values in the context of HRI. This approach aligns with previous research, such as pommeranz2012social, which successfully used expert focus groups to explore complex themes in technology and negotiation.
@@ 4.2 Focus group activities
We conducted the focus groups, which lasted approximately two hours, using an online video conference tool; the sessions were recorded for later analysis. To facilitate collaboration, we used Miro [^3] https://miro.com/, an online platform that allows users to interact with interactive boards that can be prepared in advance. The focus groups consisted of three structured activities: (1) HRI scenario analyses (individual task), (2) values review (group task), and (3) contextual information and relationship with values (group task).
These activities, designed to leverage the dynamics of group discussion while ensuring individual reflection, are inspired by previous research in related fields. The use of scenarios and visual aids in focus groups is supported by HCI research, which demonstrates that such tools stimulate richer discussions [goodman_Gathering_Requirements]. Our methodology applies it to the HRI field as in spitale2024appropriateness, who used a scenario-based evaluation to assess whether the language generated by large language models was appropriate for a robotic coach promoting mental wellbeing.
Each session was guided by two researchers: one introduced the activities and acted as a moderator to facilitate discussion and stimulate conversation, without interrupting the discussion flow of the participants and avoiding imposing external views, while the other observed and took detailed notes. This setup was chosen to minimise researcher influence on participant interactions.
@@ 4.2.1 Activity 1: scenarios analysis
The first activity involved analysing four distinct HRI fictional value scenarios provided to the participants. A value scenario is an extension of scenario-based design that helps envision the systemic effects of new technologies [nathan2007value]. We developed our HRI value scenarios through a structured process. First, we identified the main topics for our value scenarios based on the results of the previous phase and previous research [lambert2020systematic]. Specifically, we considered healthcare (focused both on elderly care [broekens2009assistive] and autism therapy [diehl2012clinical]), education (focused on children’s learning [mubin2013review]), and home assistance [sung2008housewives] as our application contexts (matching the keywords of the search query in Phase 1). Second, we created four value scenarios, one for each identified HRI domain, using ChatGPT [^4] https://chatgpt.com/, as demonstrated by jung2023toward. We followed the same approach to draft prompts for generating these scenarios. The specific prompts used to generate the scenarios are provided in Appendix A. Third, we refined the generated value scenarios by removing any references to potential moral values to avoid biasing the scenarios. The scenarios generated are reported in Appendix B.
We provided participants with these value scenarios one by one and asked them to individually identify the moral values present in each scenario. The participants provided written responses on virtual sticky notes in Miro. This individual reflection phase was designed to ensure all participants, including less confident members, had the opportunity to articulate their thoughts before the group discussion. Following this, the moderator initiated a group discussion, encouraging participants to share and debate their perspectives. This activity lasted approximately 40 minutes, 10 minutes for each scenario.
@@ 4.2.2 Activity 2: values review
The second activity involved presenting participants with the results from Phase 1, which included a list of topics and values identified and extracted from the literature. To make these HRI moral values easier to visualise, we organised them in both a table (see Table 2) and a wheel (as shown in Figure 5). We first asked participants to take their time reviewing the values, then encouraged them to collaboratively discuss and reflect on them. Their opinions were transcribed in real time in Miro by one of the researchers to support the discussion. This activity lasted approximately 20 minutes.
@@ 4.2.3 Activity 3: contextual information and relationship with values
The final activity involved discussing the relationship between the identified values and the contextual information. Our literature review revealed that most values are highly context-dependent in line with previous literature [de2016ethical]. Therefore, we decided to prioritise context as the main lens for framing these values more effectively in any HRI scenario. To this end, we have designed an activity where we asked participants to discuss the relationship between the values they have reviewed and the contextual information of HRI scenarios taking inspiration from theories in Contextual Design [holtzblatt1997contextual]. We specifically prompted them using the following contextual questions (as depicted in Figure 4): (1) Who are the users and robots? What are their demographics (age, gender, education, etc.)? What are their roles and responsibilities?; (2) What are the users’ goals? What are the primary tasks they need to accomplish? What are their motivations for using the robot?; (3) Where is the robot being used? Is it being used in a specific physical location (e.g., home, office, outdoors)? What environmental factors (lighting, noise, space) influence its use?; (4) When (how long) is the robot being used? Is there a specific time of day, week, or year when the robot is used most? How does the timing affect the user’s interaction with the robot?; (5) Why is the robot being used? What problems or needs does the robot address? What are the user’s expectations from the robot?; (6) How is the robot being used? What steps do users take to accomplish their tasks? What tools or devices do they use in conjunction with the robot?; (7) What are the social and cultural contexts? Are there social norms or cultural factors that influence how the robot is used? How do interactions with other people (colleagues, family, friends) affect usage?; (8) What are the constraints and limitations? Are there any physical, cognitive, or technological limitations users face? What budgetary, time, or resource constraints exist? This activity lasted around one hour.
@@ 4.2.4 Data analysis
We first automatically transcribed the two focus group recordings and then manually reviewed them for accuracy. We used the thematic analysis method [braun2012thematic], with the values identified during the scoping review serving as main themes and the topics as codes.
@@ 4.3 Results: topics and values in HRI
In Activity 1, the participants were asked to identify the values in the fictional HRI scenarios presented to them, without any influence from our preliminary scoping review findings. Activity 2 revealed the gaps in the values identified by the scoping review, while Activity 3 highlighted the link between contextual information and moral values. The participants noted that each piece of contextual information could be associated with multiple moral values, though they focused on the most relevant connections during discussions. At the end of the focus group, we revised the initial value definitions from the scoping review, and based on the outputs of the focus groups, we defined 26 topics grouped under 8 values, as shown in Table 5. Each topic has a positive or negative valence and is visualised in Figure 5. For clarity, we combined the results from all activities and we report the final list in the following sections. Appendix D provides a detailed overview of the thematic analysis, including citations, quotes, codes, and themes from participants organised by activity. Our analysis reached data saturation [hennink2022sample], allowing us to confidently identify and understand the final list of values, which are described in the following paragraphs. Note that this is a simplified and reductionist framework designed to assist researchers in identifying and addressing values within the HRI scenarios they are developing and not a comprehensive tool that captures the full complexity of ethical concerns and values, cultural contexts, and stakeholders’ perspectives in the HRI realm.
@@ 4.3.1 Value: Agency
This value encompasses three topics – independence, dependence, and control – all related to the user’s physical freedom. This term, prior to the focus groups, referred to one of the topics, but it was promoted to a value following the suggestions from the focus groups participants, uniforming our definition of agency with prunkl2022human: “A person is able to act on the beliefs and values they hold. This implies that they have meaningful options available to them, allowing them to make choices that are of practical import to their life”. As it emerged during the focus groups, agency differs from autonomy as the latter is used in this work to encompass topics which do not possess physical connotations. The related topics are defined as follows.
@@
The use of robots in daily life can significantly enhance a user’s independence, by reducing their physical dependence on other people, contributing to a sense of personal agency. This is especially evident in the case of elderly individuals or people with physical disabilities who require constant assistance. We included in this topic the topics about ease of use of the robot appliance, which should not require the user to rely on someone else to use the robot, as emerged during the scoping review [street2022older, vandemeulebroucke2020ethics].
@@
The previous topic can also be viewed with negative connotations [cappuccio2021can, vcaic2018service]. Indeed, when the subject’s independence relies on robots instead of other humans, the subject becomes dependent on the robots themselves. For example, the focus groups’ participants mentioned that this could lead to a deterioration of various skills, such as cooking or reasoning. One of the participants particularly questioned if deskilling might be offset by the development of new, possibly superior, skills gained through interaction with the robot.
@@
To ensure the user’s agency, robots should be designed to never physically constrain or force the user into actions they do not wish to take. During the scoping review, this topic was initially categorised under the value empowerment (e.g., [vandemeulebroucke2020ethics]). However, experts suggested removing empowerment as a separate value and instead placing control under the value agency. The required level of control can vary greatly depending on the context in which a robot is used, as noted in the focus group discussions. For example, a robot assisting in physical rehabilitation should follow the user’s pace and respond to their commands rather than pushing them too hard or acting autonomously in a way that feels coercive.
@@ 4.3.2 Value: Connectedness
This value covers the social aspects involved in human-robot interactions, especially in older adults [register2010quality]. The focus group participants suggested that this moral principle is linked to whether the robot is replacing a human role, enhancing connections, or enabling new activities. They emphasised that the robot should not replace humans, but support them.
@@
Robots have the potential to provide valuable companionship for individuals who feel isolated or lonely, e.g., [jung2023enjoy, huber2016ethical]. A companion robot can engage in conversations, respond to emotional cues, and even provide comfort just through its presence. For example, elderly individuals who live alone might find a robotic companion to be a source of emotional support, helping to alleviate feelings of loneliness.
@@
Beyond providing direct companionship, robots can also facilitate human connections, as highlighted by amirabdollahian2013accompany. For example, robots can help the elderly stay in touch with their families. This can also happen when they serve as conversation starters or mediators in social settings during group events. In other instances, they achieve the same objective by connecting the user with friends and family through video calls.
@@
On the negative side, while robots can enhance social connectedness, there is also a risk that their use might become a substitute for human interaction as emerged during the scoping review (e.g., [yeung2023living, street2022older]). This could happen when a person begins to prefer the company of their robot over that of other people. The focus group participants provided several examples of this. For instance, family members might feel less compelled to check on their elderly relatives, knowing that a robot is providing companionship. Similarly, teachers may not engage effectively with students or provide meaningful feedback if a robot takes over all teaching activities.
@@ 4.3.3 Value: Privacy
The value of privacy [finn2013seven] is the one encompassing the largest amount of topics as highlighted in both the scoping review (e.g., in [smakman2021moral]) and the focus groups. It includes – among others – the right to be informed, to access and share the data collected, and the issues related to being continuously under the watchful eye of a robot. The focus group participants have also highlighted that this value is strongly related to where, how and in which context the robot is used. Using a robot at home is completely different from using it in a public space. The experts emphasised that privacy should be maintained even when users do not want others to know they are using the robot.
@@
From the scoping review emerged the importance of ensuring that the user has access to the data collected by the robot [draper2017ethical]. Additionally, it was noted that sharing recorded information with selected individuals can be beneficial, for instance, sharing biometric data with a chosen physician. The focus group discussions emphasised that data protection is particularly crucial when dealing with vulnerable populations, such as children and the elderly. It is also vital to clearly disclose upfront who will have access to the collected data.
@@
One of the potential downsides of robots is the risk of invading the privacy of bystanders who have not consented to being monitored, e.g., [hungtechnological]. This is a concern especially in hospitals, as the robot is surrounded by vulnerable subjects in delicate settings. For example, the focus group participants questioned whether patients in nursing homes can truly give informed consent to data collection and use, given their potentially limited understanding.
@@
The constant presence of a robot in a user’s personal space can lead to feelings of discomfort and a sense of being constantly observed that can violate the value of intimacy, e.g., [aldinhas2017robots]. Furthermore, this can lead to an unconscious change in the users’ actions, behaving in ways they think will be acceptable to potential observers, rather than acting naturally. This resembles the effects of the panopticon [foucault1977discipline], as noted by one of the focus group participants. Even when the robot is not actively monitoring the environment, its presence alone is enough to influence the user’s behaviour.
@@
This topic covers all those concerns related to the excessive quantities of data collected by robotic devices (e.g., in [ostrowski2022mixed, huber2016ethical]). The continuous and aggressive collection of personal data for profiling risks misuse by both collecting companies and malicious actors. The focus groups participants expressed concerns that data from elderly care centres could be shared with insurance companies or sold to third parties without benefitting the user. These concerns are heightened for vulnerable groups like children, where parents or guardians should manage data consent. Additionally, the context of the application, such as in-home settings with less stringent regulations, can increase potential risks and harms, such as misuse by domestic abusers [winkle2024anticipating]. Other concerns, emerged during the scoping review, touch on the safe transmission and storage of this data.
@@
Transparency about the robot’s capabilities, especially regarding data collection and monitoring, emerges with positive connotations in the scoping review. Users should be fully informed about what the robot is capable of, and what data is being shared, as highlighted by focus group participants. This can be tricky to implement when transparency would reduce the effectiveness of the robot itself, in the case of personal disclosure for instance.